# Noise Contrastive Estimation-based Matching Framework for Low-resource Security Attack Pattern Recognition

### v1

@@ 1 Introduction and Background
Cyber Threat Intelligence (CTI), an essential pillar of cybersecurity, involves collecting and analyzing information on cyber threats, including threat actors, their campaigns, and malware, helping timely counterintelligence and defending efforts. Textual threat reports or blogs are considered a crucial source of CTI, where security vendors diligently investigate and promptly detail intricate attacks. A key and intricate sub-task in extracting CTI from these textual sources involves the identification of Tactics, Techniques, and Procedures (TTP) of the threat actors, i.e. comprehending descriptions of low-level, complex threat actions and connecting them to standardized attack patterns. One of the popular standard knowledge frameworks widely adopted in the CTI community is MITRE ATT&CK Storm et al. (2020). Within this framework, a technique represents a specific method used to achieve an objective, with its corresponding tactics and sub-techniques covering broader strategies and variations. Figure 1 illustrates an example of a text in a threat report, of which indicating two attack patterns, among others, i.e., (1) the use of a malicious email attachment to take control of a victim’s system (T1566 111attack.mitre.org/techniques/T1566), and (2) encrypting data on the victim’s system for ransom demands. (T1486 222attack.mitre.org/techniques/T1486).
@@
As of 2023, there are over 600 techniques, together with 14 high-level tactics described in MITRE ATT&CK. In its ontology, a technique is associated to at least one tactic (e.g., the technique “Hijack Execution Flow” is listed under three distinct tactics: Persistence, Privilege Escalation and Defense Evasion) and may have several sub-techniques. Mining techniques from CTI reports poses significant challenges due to several factors. Firstly, the large number of techniques, coupled with their diverse nature, intricate inter-dependencies, and hierarchical structure, renders the task complex and laborious. Secondly, the analysis of CTI reports necessitates the expertise of security professionals. The reports focus on delineating low-level threat actions rather than explicitly mentioning the associated techniques and tactics. Consequently, extracting relevant techniques and tactics from these reports requires diligent inference by the reader. Employing an automated approach to TTP mapping presents inherent challenges. One major hurdle is the low-resource nature of the task, due to the limited availability of labeled data and the extensive label space. Moreover, the presence of long-tail infrequent TTPs adds complexity to the learning process.
@@ 2 Related Work
TTP Mapping and CTI Extraction Several works target TTP mapping on the document level.  Husari et al. (2017) used a probabilistic relevance framework (Okapi BM25) to quantify the similarity between bag-of-words representations of TTPs and the target text. However, this approach is limited to the oversimplified vocabulary of threat actions within an ad-hoc ontology. Ayoade et al. (2018); Niakanlahiji et al. (2018); Legoy et al. (2020) used a TF-IDF-based document representation and leveraged classical (i.e., tree-based, margin-based) ML for (multi-label) classification. Li et al. (2019) used latent semantic analysis to extract topics from target articles, and compared the topic vectors with the TF-IDF vectors of ATT&CK description pages to obtain cosine similarity. They used the similarity vectors with Naïve Bayes and decision trees to classify TTPs. However, the choice of document-level granularity introduces additional unneeded complexity of long-form text representation. Recent works leverage transformers for sentence-level text representation learning Orbinato et al. (2022); You et al. (2022), using the encoded representation in the multi-class classification setting. However, with limited available data, they restrict the task to only a small number of TTPs.
@@ 3 Preliminaries and Problem Setup
[…]
@@
Problem Reformulation. Given the training data $\mathsf{X}\in\mathbb{R}^{n\times\mathcal{|D|}}$, and $\mathsf{Y}\in\mathbb{R}^{\mathsf{|L|}\times\mathcal{|D|}}$, with $y\in\mathsf{Y}$ derived from the TTP textual profile, and $|\mathsf{L}|\ll\infty$ along with a set of supervisions $\{x\mapsto y\}^{n}=\{0,1\}^{n}$, such as $x\in\mathsf{X}$ and $y\in\mathsf{Y}$, our target is to learn matching-based scoring functions $\mathsf{g}_{\theta}(x,y)$ that model the relationship between $x$ and $y$ within the same feature space, aiming for $\mathsf{g}_{\theta}(x,y)\approx\{x\mapsto y\}^{n}$. The use of the textual profile inherently eliminates the need for an projection function $\pi$, as it directly aligns the discrete categorical representation Y with the same continuous space as X. In the context of cross-entropy loss, $p_{\theta}(y|x)$ is now linked to $p_{\theta}(x\mapsto y|x,y)$.
@@ 4.1 Matching Network
The architecture of our matching network is built upon the dual-encoder framework, which typically employs a Siamese network. This shared network is used for learning the representations of both the target text segment and the TTP textual profile. As depicted in Figure 2, at a high level, our network comprises an embedding component and an alignment component. Each includes specific layers aimed at enhancing the connectivity between the two sub-network sides. Finally, the two sides are merged (by i.e., a dot product) to output a (probabilistic) matching score. We detail the architectural choice for our matching network below.
@@
Encoder. The encoder has two modes: (1) scratch and (2) scratch with a pre-trained transformer (i.e., SecBERT) combined. Scratch indicates that the token embedding are learnt (with the embedding layer). We then apply a simple CNN on top of the embedding layer. With scratch alone, a specialized tokenizer (that respect CTI entities, e.g., URL, vulnerability identifier..) is used. While using together with the transformer, the tokenizer of the transformer is used. For (2), we simply stack the encoded vectors from the two sources together.
Alignment Network. Formally, given the input representation of the text-TTP pair as $x_{t}=(\hat{a}_{1},\ldots,\hat{a}_{l})$ and $y_{ttp}=(\hat{b}_{1},\ldots,\hat{b}_{l})$, the unnormalized attention weights are decomposed into: $e_{ij}=W^{align}(\hat{a_{i}})\cdot W^{align}(\hat{b_{j}})$, whereas $W^{align}$ is a trainable projection matrix, $\cdot$ is the dot product. Then, we derive the normalized weights for each token $a_{i}$ and $b_{j}$, and achieve the corresponding alignment features.
Similar to  Yang et al. (2019), we use the block-based residual architecture with skip connections. Our block consists of the encoder, alignment and fusion layers. The fusion layer does various comparisons of local and aligned representations (i.e., the Hadamard product) and finally fuses the interaction vectors together using the concatenation operator. Then pooling, i.e., (non-) weighted average or max-pooling, is applied to attain fixed-length vector representations.
4.2 Learning by Contrasting
Our efficient learning method aims to circumvent the computational complexities that arise in the large label space, whether in the proper multi-label setting or its reduced multi-class version. The new learning paradigm is shifted from multi-label classification to the so-called dynamic label-informed text matching, in which negative labels are drawn dynamically at every step. The ranker, acting as a simultaneous matcher, strategically assigns higher probabilities to positive pairs and lower probabilities to negative pairs. Finally, the top-k positive pairs are selected based on a cut-off threshold. We detail our learning mechanism below.
Partial-ranking-based NCE. The general idea of NCE in our scenario is to avoid an exhaustive ranking (or partitioning) in the large label space, i.e., in the vanilla multi-label classification setting. Instead, a matching-based classifier, $p((x\mapsto y)|x,y)$, is trained to differentiate between samples from the true distribution and a noise distribution, $q(y)$, and inherently approximate the underlying ranking function. By utilizing Monte Carlo sampling, the NCE loss can be formulated as follows:
@@ 4.2 Learning by Contrasting
While the NCE loss in Equation 2 is calculated by learning $p((x\mapsto y)|x,y)$ for every data point (so-called local), we opt for a ranking setting where data points in the same batch compete in a contrastive setting. One way of achieving this is to use the mutual information, as utilized in InfoNCE Oord et al. (2018), to quantify the distance between the prediction and (k-sized) label distributions. The ranking NCE loss is then defined as:
@@
whereas, $g_{\theta}(x,y)$ is the matching function. Consequently, minimizing the loss promotes simultaneously a lower $g_{\theta}$ for negative pairs and a higher $g_{\theta}$ for positive pairs. The scaling factor $\gamma$, which is absent in InfoNCE, is introduced to account for the need to reduce the impact of the considerably larger portion of negative samples. This adjustment aims to emphasize the top-k partial ranking, where it is assumed that the positive samples are concentrated in the distribution. Subsequently, when $\gamma$ presents, the loss is denoted as $\alpha$-balanced NCE.
@@ 4.3 Sampling Strategies
Corpus-level negative sampling. Due to memory constraints, the conventional negative sampling method is often applied in-batch Yih et al. (2011); Gillick et al. (2019). One limitation of the in-batch sampling is the number of negative samples are bounded to the batch size. Whereas, the corpus-level sampling provides a broader context for negative sampling, inherently leading to a more diverse set of negative examples. In our low-resource context, the diversified negative samples is extremely useful in enhancing the discriminative power of the dataset, that is likely not evident within a single batch. In effect, we assume that a larger part of the TTP corpus is irrelevant to the positive paired sample. We also assume that noisy samples will inherently be canceled out while learning signals remain in our training paradigm Rolnick et al. (2017). While being simple, the policy augments our dataset with a substantial supervision signal stemming from negative samples. We explain the details of our sampling policies below.
@@ 5.1 Datasets
Procedure+.666 The datasets will be publicly shared at github.com/TTP-Mapping to foster further research. Procedures: collected from ATT&CK, where techniques have associated manually curated procedure examples footnote 3. Each example is a one-sentence expert-written summary of the implementation of a technique in real-world attacks. Derived Procedures: complements an example with a text that aligns to threat report writing style. We look for evidential paragraphs in the references where the summary example is assumedly derived from, using a per-document search engine.
Expert.footnote 6 Our purposefully crafted dataset closely emulates real-world scenarios, providing an practical setting for TTP extraction. Unlike sentence-focused datasets, ours covers entire paragraphs, thus the annotations are inherently multi-label in nature. Annotated by 5 CTI experts using an in-house tool, our dataset triples text length and increases average labels per sample by approximately 60-80% compared to TRAM (see Table1).
@@
Table 2: Results of all models on 3 datasets. Procedures+ denotes the combined procedure examples datasets. Bold denotes best while underscore signifies second-best performance. Indented ()(-)( - ) denotes training without the specific option wrt. the preceding model. Ideal R@$1$ on the Expert dataset is 0.504. ${}^{\mathcal{T}}$ uses pre-trained SecBERT.
@@
Baseline
TTPDrill (BM25) .230 .227 .118 .232 .250 .212 .118 .205 .222 .037 .008 .139
[…]
Dynamic ${}^{\mathcal{T}}$339 .336 .277 .432 .286 .253 .277 . 402 .449 .111 .252 .525
XMTC
eXtremeText (Sigmoid) .557 .547 .371 .624 .632 .594 .425 .729 .407 .174 .279 .485
@@
XR-${}^{\mathcal{T}}$502 .494 .304 .548 .540 .515 .334 .595 .389 .149 .239 .453
Ours
[…]
   ()(-)( - ) auxiliary .604 .584 .433 .719 .712 .601 .521 .816 .693 .177 .442 .773
   ()(-)( - ) Transformers .646 .601 .357 .772 .642 .543 .547 .785 .700 .173 .430 .766
[…]
@@
Baseline
TTPDrill (BM25) .294 .290 .152 .297 .281 .271 .161 .295 .197 .096 .096 .279
[…]
Dynamic ${}^{\mathcal{T}}$449 .447 .408 .539 .404 .353 .382 .513 .559 .166 .344 .631
XMTC
eXtremeText (Sigmoid) .659 .649 .426 .713 .742 .704 .494 .793 .439 .212 .333 .521
@@
XR-${}^{\mathcal{T}}$683 .673 .416 .747 .801 .750 .488 .856 .554 .245 .405 .633
Ours
[…]
   ()(-)( - ) auxiliary .714 .689 .579 .791 .817 .697 .648 .88 .754 .233 .509 .816
   ()(-)( - ) Transformers .777 .733 .664 .86 .791 .683 .713 .875 .718 .226 .497 .782
[…]

### v4

@@ 1 Introduction and Background
Cyber Threat Intelligence (CTI), an essential pillar of cybersecurity, involves collecting and analyzing information on cyber threats, including threat actors, their campaigns, and malware, helping timely threat detection and defense efforts. Textual threat reports or blogs are considered a important source of CTI, where security vendors diligently investigate and promptly detail intricate attacks. A key sub-task in extracting CTI from these textual sources involves the identification of Tactics, Techniques, and Procedures (TTP) of the threat actors, i.e. comprehending descriptions of low-level, complex threat actions and connecting them to standardized attack patterns. One of the popular standard knowledge frameworks widely adopted in the CTI community is MITRE ATT&CK Storm et al. (2020). Within this framework, a technique represents a specific method used to achieve an objective, with its corresponding tactics and sub-techniques covering broader strategies and variations. Fig. 1 illustrates an example of a text in a threat report, which indicates two attack patterns, among others, i.e., (1) the use of a malicious email attachment to take control of a victim’s system (T1566 111attack.mitre.org/techniques/T1566), and (2) encrypting data on the victim’s system, presumably for ransom demands (T1486 222attack.mitre.org/techniques/T1486).
@@
As of 2024, there are over 600 techniques, together with 14 high-level tactics described in MITRE ATT&CK. In its ontology, a technique is associated to at least one tactic (e.g., the technique “Hijack Execution Flow” is listed under three distinct tactics: Persistence, Privilege Escalation and Defense Evasion) and may have several sub-techniques. Mining techniques from CTI reports poses significant challenges due to several factors. Firstly, the large number of techniques, coupled with their diverse nature, intricate inter-dependencies, and hierarchical structure, renders the task complex and laborious. Secondly, the analysis of CTI reports necessitates the expertise of security professionals. The reports focus on delineating low-level threat actions rather than explicitly mentioning the associated techniques and tactics. Consequently, extracting relevant techniques and tactics from these reports requires diligent inference by the reader. Employing an automated approach to TTP mapping presents inherent challenges. One major hurdle is the low-resource nature of the task, due to the limited availability of labeled data and the extensive label space. Moreover, the presence of long-tail infrequent TTPs adds complexity to the learning process.
@@ 2 Related Work
TTP Mapping and CTI Extraction Several works target TTP mapping on the document level.  Husari et al. (2017) used a probabilistic relevance framework (Okapi BM25) to quantify the similarity between BoW representations of TTPs and the target text. However, this approach is limited to the oversimplified vocabulary of threat actions within an ad-hoc ontology. Ayoade et al. (2018); Niakanlahiji et al. (2018) used a TF-IDF-based document representation and leveraged classical (i.e., tree-based, margin-based) ML for (multi-label) classification. Li et al. (2019) used latent semantic analysis to extract topics from target articles, and compared the topic vectors with the TF-IDF vectors of ATT&CK description pages to obtain cosine similarity. They used the similarity vectors with Naïve Bayes and decision trees to classify TTPs. However, the choice of document-level granularity introduces additional unneeded complexity of long-form text representation. Recent works leverage transformers for sentence-level text representation learning Orbinato et al. (2022); You et al. (2022), using the encoded representation in the multi-class classification setting. However, with limited available data, they restrict the task to only a small number of TTPs.
@@ 3 Preliminaries and Problem Setup
Classification: Let X and Y denote the input and label spaces, $|\textsf{Y}|<\infty$. We define a score function $\textsl{g}_{\theta}:\textsf{X}\rightarrow\textsf{Y}$. In this setting, the label space Y is categorical. Specifically, $\textsf{X}\in\mathbb{R}^{n\times m}$, whereas $\textsf{Y}\in\{0,1\}^{n\times|L|}$, with $n$ being the number of samples and $L$ being the label set.
Matching: In this setting, X and Y represent the same input space. The matching function $\textsl{g}_{\theta}:\textsf{X}\times\textsf{Y}\rightarrow\mathbb{R}$, is differentiable in $\theta\in\mathbb{R}^{|\mathcal{D}|}$, where $D$ is the parameter space. In order to cast a classification problem as a matching one, we assume there is an invertible and smooth projection function $\pi$ that transforms the discrete categorical representation Y into the same continuous space as X.
@@
Problem Reformulation. Given the training data $\mathsf{X}\in\mathbb{R}^{n\times\mathcal{|D|}}$, and $\mathsf{Y}\in\mathbb{R}^{\mathsf{|L|}\times\mathcal{|D|}}$, with $y\in\mathsf{Y}$ derived from the TTP textual profile, and $|\mathsf{L}|\ll\infty$ along with a set of supervisions $\{x\mapsto y\}^{n}=\{0,1\}^{n}$, such as $x\in\mathsf{X}$ and $y\in\mathsf{Y}$, our target is to learn matching-based scoring functions $\mathsf{g}_{\theta}(x,y)$ that model the relationship between $x$ and $y$ within the same feature space, aiming for $\mathsf{g}_{\theta}(x,y)\approx\{x\mapsto y\}^{n}$. The use of the textual profile inherently eliminates the need for a projection function $\pi$, as it directly aligns the discrete categorical representation Y with the same continuous space as X. In the context of cross-entropy loss, $p_{\theta}(y|x)$ is now linked to $p_{\theta}(x\mapsto y|x,y)$.
@@ 4.1 Matching Network
The architecture of our matching network is built upon the dual-encoder framework, which typically employs a Siamese network. This shared network is used for learning the representations of both the target text segment and the TTP textual profile. As depicted in Fig. 2, at a high level, our network comprises an embedding component and an alignment component. Each includes specific layers aimed at enhancing the connectivity between the two sub-network sides. Finally, the two sides are merged (by, i.e., a dot product) to output a (probabilistic) matching score. We detail the architectural choice for our matching network below.
@@
Encoder. The encoder has two modes: (1) scratch and (2) scratch with a pre-trained transformer (i.e., SecBERT) combined. Scratch indicates that the token embeddings are learnt (with the embedding layer). We then apply a simple CNN on top of the embedding layer. With scratch alone, a specialized tokenizer (that respects CTI entities, e.g., URL, vulnerability identifier..) is used. While using together with the transformer, the tokenizer of the transformer is used. For (2), we simply stack the encoded vectors from the two sources together.
Alignment Network. Formally, given the input representation of the text-TTP pair as $x_{t}=(\hat{a}_{1},\ldots,\hat{a}_{l})$ and $y_{ttp}=(\hat{b}_{1},\ldots,\hat{b}_{l})$, the unnormalized attention weights are decomposed into: $e_{ij}=W^{align}(\hat{a_{i}})\cdot W^{align}(\hat{b_{j}})$, whereas $W^{align}$ is a trainable projection matrix, $\cdot$ is the dot product. Then, we derive the normalized weights for each token $a_{i}$ and $b_{j}$, and achieve the corresponding alignment features. Similar to  Yang et al. (2019), we further use the block-based residual architecture with skip connections. Our block consists of the encoder, alignment and fusion layers. The fusion layer does various comparisons of local and aligned representations (i.e., the Hadamard product) and finally fuses the interaction vectors together using the concatenation operator. Then pooling, i.e., (non-) weighted average or max-pooling, is applied to attain fixed-length vector representations.
4.2 Learning to Match and Contrast
Our efficient learning method aims to circumvent the computational complexities that arise in the large label space, whether in the proper multi-label setting or its reduced multi-class version. The new learning paradigm is shifted from multi-label classification to the so-called dynamic label-informed text matching, in which negative labels are drawn dynamically at every step. The ranker, acting as a simultaneous matcher, strategically assigns higher probabilities to positive pairs and lower probabilities to negative pairs. Finally, the top-$n$ positive pairs are selected based on a cut-off threshold. We detail our learning mechanism below.
Partial-ranking-based NCE. The general idea of NCE in our scenario is to avoid an exhaustive ranking (or partitioning) in the large label space, i.e., in the vanilla multi-label classification setting. Instead, a matching-based classifier, $p((x\mapsto y)|x,y)$, is trained to differentiate between samples from the true distribution and a noise distribution, $q(y)$, and inherently approximate the underlying ranking function. By utilizing Monte Carlo sampling, the NCE loss is formulated as follows:
@@ 4.2 Learning to Match and Contrast
While the NCE loss in Equation 2 is calculated by learning $p((x\mapsto y)|x,y)$ for every data point (so-called local), we opt for a ranking setting where data points in the same batch compete in a contrastive setting. One way of achieving this is to leverage the mutual information $\mathcal{I}$, as utilized in InfoNCE Oord et al. (2018), to encourage informative representations for the positive samples $\mathcal{I}(z(x,y);z(x,y^{(+)}))$ (assuming multi-label setting) and contrast them with negative ones $\mathcal{I}(z(x,y);z(x,y^{(-)}))$. The ranking NCE loss is then defined as:
@@
whereas, $g_{\theta}(x,y)$ is the matching function. Consequently, minimizing the loss promotes simultaneously a lower $g_{\theta}$ for negative pairs and a higher $g_{\theta}$ for positive pairs. The scaling factor $\gamma$, which is absent in InfoNCE, is introduced to account for the need to reduce the impact of the considerably larger portion of negative samples. This adjustment aims to emphasize the top-$n$ partial ranking, where it is assumed that the positive samples are concentrated in the distribution. Subsequently, with $\gamma$, the loss is denoted as $\alpha$-balanced NCE.
@@ 4.3 Sampling Strategies
Corpus-level negative sampling. Due to memory constraints, the conventional negative sampling method is often applied in-batch Yih et al. (2011); Gillick et al. (2019). One limitation of the in-batch sampling is the number of negative samples are bounded to the batch size. Whereas, the corpus-level sampling provides a broader context for negative sampling, inherently leading to a more diverse set of negative examples. In our low-resource context, the diversified negative samples are extremely useful in enhancing the discriminative power of the dataset, that is likely not evident within a single batch. In effect, we assume that a larger part of the TTP corpus is irrelevant to the positive paired sample. We also assume that noisy samples will inherently be canceled out while learning signals remain in our training paradigm Rolnick et al. (2017). While being simple, the policy augments our dataset with a substantial supervision signal stemming from negative samples. We explain the details of our sampling policies below.
@@ 5.1 Datasets
Procedure+.666 The datasets are publicly shared at github.com/TTP-Mapping to foster further research. Procedures: collected from ATT&CK, where techniques have associated manually curated procedure examplesfootnote 3. Each example is a one-sentence expert-written summary of the implementation of a technique in real-world attacks. Derived Procedures: complements an example with a text that aligns to threat report writing style. We look for evidential paragraphs in the references where the summary example is assumedly derived from, using a per-document search engine.
Expert.footnote 6 Our purposefully crafted dataset closely emulates real-world scenarios, providing a practical setting for TTP extraction. Unlike sentence-focused datasets, ours covers entire paragraphs, thus the annotations are inherently multi-label in nature. Annotated by 5 CTI experts using an in-house tool, our dataset triples text length and increases average labels per sample by approximately 60-80% compared to TRAM (see Table1).
@@
Table 2: Results of all models on 3 datasets. Procedures+ denotes the combined procedure examples datasets. Bold denotes best while underscore signifies second-best performance. Indented (w/o) denotes training without the specific option wrt. the preceding model. Ideal R@$1$ on the Expert dataset is 0.504. T uses pre-trained SecBERT.
@@
Baseline TTPDrill (BM25) .230 .227 .118 .232 .250 .212 .118 .205 .222 .037 .008 .139
Binary RelevanceT .206 .579 .193 .579 .236 .594 .209 .594 .189 .256 .085 .256
Dynamic Triplet-lossT .339 .336 .277 .432 .286 .253 .277 . 402 .449 .111 .252 .525
XMTC eXtremeText (Sigmoid) .557 .547 .371 .624 .632 .594 .425 .729 .407 .174 .279 .485
@@
XR-TRANSFORMERT .502 .494 .304 .548 .540 .515 .334 .595 .389 .149 .239 .453
Ours InfoNCET .672 .639 .442 .758 .697 .577 .516 .799 .702 .175 .432 .768
@limit-from@@-@ -balancedT .760 .720 .489 .837 .765 .646 .546 .856 .693 .169 .400 .762
     w/o auxiliary .604 .584 .433 .719 .712 .601 .521 .816 .693 .177 .442 .773
     w/o Transformers .646 .601 .357 .772 .642 .543 .547 .785 .700 .173 .430 .766
AsymmetricT .757 .718 .493 .838 .770 .658 .555 .864 .731 .182 .399 .789
@@
Baseline TTPDrill (BM25) .294 .290 .152 .297 .281 .271 .161 .295 .197 .096 .096 .279
Binary RelevanceT .409 .655 .285 .655 .399 .647 .279 .647 .167 .295 .117 .295
Dynamic Triplet-lossT .449 .447 .408 .539 .404 .353 .382 .513 .559 .166 .344 .631
XMTC eXtremeText (Sigmoid) .659 .649 .426 .713 .742 .704 .494 .793 .439 .212 .333 .521
@@
XR-TRANSFORMERT .683 .673 .416 .747 .801 .750 .488 .856 .554 .245 .405 .633
Ours InfoNCET .759 .727 .624 .823 .819 .696 .668 .876 .741 .228 .515 .871
@limit-from@@-@ -balancedT .843 .806 .666 .892 .889 .778 .711 .927 .731 .224 .491 .789
     w/o auxiliary .714 .689 .579 .791 .817 .697 .648 .88 .754 .233 .509 .816
     w/o Transformers .777 .733 .664 .86 .791 .683 .713 .875 .718 .226 .497 .782
AsymmetricT .841 .806 .677 .892 .903 .789 .726 .938 .745 .236 .483 .802